OpenStudy (anonymous):

On the graph of the equation 4x + 5y = 20, what is the value of the y-intercept? -5 -4 4 5

OpenStudy (aum):

To find the y-intercept, set x = 0 in the equation and solve for y.

OpenStudy (anonymous):

0x + 5y = 20, I don't just divide by 5, thats too simple.

OpenStudy (anonymous):

Unless I do

OpenStudy (anonymous):

I feel like I do but it just seems too easy.

jimthompson5910 (jim_thompson5910):

more like 4x + 5y = 20 4*0 + 5y = 20 0 + 5y = 20 5y = 20 y = ???

OpenStudy (anonymous):

Wow these are easy, I have to check up on this.

OpenStudy (anonymous):

4, thank you.

jimthompson5910 (jim_thompson5910):

yep, it's not too bad